*27 The potential difference across the I-00 μF capacitor in Fig. 18-58 is 10.0 V. Find (a) the equivalent capacitance between A and B: (b) the potential difference between C and D 1.00 F2.00 pFF 2.00 uk: 3 (H) İlF 7 00 Fig. 18-58
Answer #1

a) for the equivalent capacitance

as 1 uF and 2 uF and 2 uF in series

1/C1 = 1/1 + 1/2 + 1/2

C1 = 0.50 uF

Now,C1 and 3 uF in parallel

C2 = 0.50 + 3 = 3.5 uF

NOw, for C2 and uF in series

1/Ceq = 1/3.5 + 1/7

Ceq = 2.33 uF

the equivalent capacitance is 2.33 uF

b) Now, as the potential across

let the potential across C and D

VCD * C1 = 10 * 1

VCD = 10 * 1/0.50 = 20 V

the potential difference between C and D is 20 V